OpenStudy (anonymous):

A store mixes Brazillian coffee worth $12 per kg and Venezuelan coffee worth $16 per kg. The mixture is to sell for $14 per kg. Find how much of each should be used to make a 528-kg mixture.

Directrix (directrix):

Let B = the number of kg of Brazillian coffee Let V = the number of kg of Venezuelan coffee B + V = 528 12B is the amount of money invested in B coffee 16V is the amount of money invested in V coffee 12B + 16 V = 528 ( 14) = 7 392 528 ( 14) is the total money value of 528 kg at $14 each Solving simultaneously, B +V = 528 12 B + 16 V - 7 392 Solution by substitution: B = 528 - V 12 ( 528 - V ) + 16 V = 7 392 6 336 - 12 V + 16 V = 7 392 4V = 1 056 V = 264 kg Back substituting for V in B = 528 - V yields B = 528 - 264 = 264 kg. Therefore, B = V = 264 kg. Here's hoping somebody checks this. :)